A 900 kg car has a net force of 2000 N acting on it. What is the acceleration of the car? Express your answer in meters per second squared and round to one decimal place
Answer #1

2.2

Force=mass×acceleration

Therefore, acelararation a=F÷m

a=2000÷900=2.222m/s²

In one decimal place acelararation

a= 2.2 m/s²